# Perspective: Interactions mediated by atoms, photons, electrons, and excitons

### v1

@@ VIII.1 Fermi polaron-polariton interactions
In a recent experiment using a semiconductor monolayer of MoSe2, the presence of a 2DEG was found to have strong effects on the interaction between polaron-polaritons Tan et al. (2020). A layer of graphene allowed for the tuning of the electron density and the setup was embedded in a high-finesse optical cavity consisting of a distributed-Bragg-reflector to increase the photon-exciton coupling. First a pump beam created excitons, which due to interactions with electrons in the 2DEG involving a trion, i.e. a bound state between an exciton and an itinerant electron, formed Fermi polaron-polaritons. A broad probe beam was then applied after a delay time to create and measure the energy of polaron-polaritons in the presence of the polaron-polaritons already created by the pump beam. As seen in Fig. 11 (top), a transient blue shift of the measured polaron-polariton energy was observed for certain delay times between the pump and probe pulses. This blue shift was interpreted as a repulsive interaction between the polariton-polaritons, which was observed to increase significantly when a 2DEG was present as shown in Fig. 11 (bottom). Here the polaron-polariton interactions (purple dots) in the presence of the 2DEG are $\sim 50$ times larger than the bare exciton-polariton interactions (orange dots). Intriguingly, the observed repulsive nature of the polariton-polariton interaction is at odds with the results in Sec. II, which predict the interaction mediated by the electrons to be attractive since the excitons to a good approximation can be regarded as point bosons.
@@
Using multi-dimensional coherent spectroscopy on a monolayer WS2 Muir et al. (2022), interactions between excitons immersed in a 2DEG were studied at time scales shorter than the typical formation time of Fermi polarons. The observed interaction effects were attributed to phase-filling effects where excitons compete for the same electrons to form polarons. In addition, a bipolaron involving excitons in different valleys with a rather large binding energy was found.
@@ IX Mediated collective phases
Exciton-polariton mediated superconductivity.- The experimental realization of exciton-polariton BECs (or coherent states) has motivated several investigations concerning how they can mediate an attractive interaction between electrons thereby inducing superconductivity in a solid state setting. Contrary to atomic gases where the interaction between the fermions can be suppressed, the mediated interaction between electrons has to overcome their Coulomb repulsion. For low condensate densities, the net interaction is typically dominated by Coulomb repulsion, and electron pairing cannot take place. At high densities however, the attractive mediated interaction becomes strong enough to support Cooper pairs and superconductivity. The first proposals studied a 2DEG sandwiched between quantum wells in which a condensate of exctiton-polaritons can mediate an attractive interaction Laussy et al. (2010, 2012); Cotleţ et al. (2016). With the arrival of TMDs, Bogoliubov mode mediated superconductivity recieved renewed attention Skopelitis et al. (2018); Sun et al. (2021a); Yang et al. (2021). This includes the development of a strong-coupling theory of condensate-mediated superconductivity Sun et al. (2021b), and the demonstration of topological superconductivity with a high critical temperature due to the light mass of the polaritons forming the condensate reducing retardation effect Julku et al. (2022). Topological $p$-wave as well as $s$-wave superconductivity mediated by excitons interacting with electrons via a trion state was analysed in Refs. Zerba et al. (2023); von Milczewski et al. (2023). Bound states of two photons have also been predicted to emerge as a consequence of an attractive interaction mediated by a polariton condensate Camacho-Guardian et al. (2021).

### v3

@@ VIII.1 Fermi polaron-polariton interactions
In a recent experiment using a semiconductor monolayer of MoSe2, the presence of a 2DEG was found to have strong effects on the interaction between polaron-polaritons Tan et al. (2020). A layer of graphene allowed for the tuning of the electron density and the setup was embedded in a high-finesse optical cavity consisting of a distributed-Bragg-reflector to increase the photon-exciton coupling. First, a narrow pump beam tuned to the lower-polariton energy created polaritons, which, due to their excitonic component, interacted with electrons in the 2DEG. The presence of a trion state, i.e., a bound state between an exciton and an itinerant electron, led to an indirect coupling between the polaritons and the trion, forming strongly interacting polaron-polaritons Sidler et al. (2016). A broad probe beam was then applied after a delay time to create and measure the energy of polaron-polaritons in the presence of the polaron-polaritons already created by the pump beam. As seen in Fig. 11 (top), a transient blue shift of the measured polaron-polariton energy was observed for certain delay times between the pump and probe pulses. This blue shift was interpreted as a repulsive interaction between the polariton-polaritons, which was observed to increase significantly when a 2DEG was present as shown in Fig. 11 (bottom). Here the polaron-polariton interactions (purple dots) in the presence of the 2DEG are $\sim 50$ times larger than the bare exciton-polariton interactions (orange dots). Intriguingly, the observed repulsive nature of the polariton-polariton interaction is at odds with the results in Sec. II, which predict the interaction mediated by the electrons to be attractive since the excitons to a good approximation can be regarded as point bosons.
@@
Using multi-dimensional coherent spectroscopy on a monolayer WS2 Muir et al. (2022), interactions between excitons immersed in a 2DEG were studied at time scales shorter than the typical formation time of Fermi polarons. The observed interaction effects were attributed to phase-filling effects where excitons compete for the same electrons to form polarons. In addition, in Ref. Muir et al. (2022) a bipolaron involving excitons in different valleys with a rather large binding energy was found.
@@ IX Mediated collective phases
Exciton-polariton mediated superconductivity.- The experimental realization of exciton-polariton BECs (or coherent states) has motivated several investigations concerning how they can mediate an attractive interaction between electrons thereby inducing superconductivity in a solid state setting. Contrary to atomic gases where the interaction between the fermions can be suppressed, the mediated interaction between electrons has to overcome their Coulomb repulsion. For low condensate densities, the net interaction is typically dominated by Coulomb repulsion, and electron pairing cannot take place. At high densities however, the attractive mediated interaction becomes strong enough to support Cooper pairs and superconductivity. The first proposals studied a 2DEG sandwiched between quantum wells in which a condensate of exctiton-polaritons can mediate an attractive interaction Laussy et al. (2010, 2012); Cotleţ et al. (2016). With the arrival of TMDs, Bogoliubov mode mediated superconductivity recieved renewed attention Skopelitis et al. (2018); Sun et al. (2021a); Yang et al. (2021). This includes the development of a strong-coupling theory of condensate-mediated superconductivity Sun et al. (2021b), and the demonstration of topological superconductivity with a high critical temperature due to the light mass of the polaritons forming the condensate reducing retardation effect Julku et al. (2022). Topological $p$-wave as well as $s$-wave superconductivity mediated by excitons interacting with electrons via a trion state was analysed in Refs. Zerba et al. (2023); von Milczewski et al. (2023). Bound states of two photons have also been predicted to emerge as a consequence of an attractive interaction mediated by a polariton condensate Camacho-Guardian et al. (2021). Recently, spinor BECs have been proposed to enhance superconductivity in TMD materials and atomic systems Bighin et al. (2022).